a man travels in a motor speed bangka at the rate of 15 kph from his barrio to the city and comes back at the rate of 12 kph. If his total time of travel is 3 hours, find the distance between his barrio to the city.
Answer by lwsshak3(11628)

a man travels in a motor speed bangka at the rate of 15 kph from his barrio to the city and comes back at the rate of 12 kph. If his total time of travel is 3 hours, find the distance between his barrio to the city.
***
let x=travel time from barrio to city
3-x=travel time from city to barrio
distance=travel time*speed
15x=12(3-x)=36-12x
27x=36
x=36/27=4/3 hrs
Distance between barrio and city=15x=15*4/3=20 km